The vertex of the parabola below is at the point (4, 6.5)
and the point (5, 4.5) is on the parabola.
What is the equation of the parabola?

a. y=-2.5(x+2)^2+4.5
b. x=-2(y-4)^2+6.5
c. y=-2(x-4)^2+6.5
d. y=-2(x-4)^2+6.5

Answer by lwsshak3(11628)

The vertex of the parabola below is at the point (4, 6.5)
and the point (5, 4.5) is on the parabola.
What is the equation of the parabola?
**
Standard form of equation for a parabola: y=A(x-h)^2+k, (h,k)=(x,y) coordinates of the vertex.
y=A(x-4)^2+6.5
Solving for A using coordinates of given point (5,4.5)
4.5=A(5-4)+6.5
4.5=A(1)+6.5
A=-2
Equation: y=-2(x-4)^2+6.5
answers are c and d which are duplicates.